OpenStudy (anonymous):

Sanford can assemble a computer in 60 minutes. Colleen can assemble a computer in 40 minutes. If they work together, how many minutes does it take them to assemble a computer?

OpenStudy (welshfella):

1/40 + 1/60 = 1/x where x is the number of minutes its takes for them to complete the work together

OpenStudy (anonymous):

100 is not an answer

OpenStudy (welshfella):

each fraction on the left is the part of the work each completes in 1 minute The sum of these 2 fractions = 1/x - the fraction they complete when working togethher

OpenStudy (welshfella):

I know because x = 100 is not a solution of the equation

OpenStudy (anonymous):

20?

OpenStudy (welshfella):

To solve this equation multiply each term by the LCD of the 3 fractions which is 120x

OpenStudy (welshfella):

Its not 20

OpenStudy (anonymous):

-.-

OpenStudy (welshfella):

try and solve the equation

OpenStudy (anonymous):

100/120=5???

OpenStudy (welshfella):

No multiply each term by 120x